OpenStudy (anonymous):

Find all points of intersection algebraically. x - y - 1 = 0 y = x^2 - x - 1

OpenStudy (anonymous):

x - x^2 - x - 1 - 1 = 0

OpenStudy (anonymous):

-x^2 - 2 = 0

OpenStudy (anonymous):

What do I do from there?

OpenStudy (michele_laino):

I solve the first equation for y, and I get this: \[y = x - 1\] then I replace y with x-1 into the second equation, so I can write: \[x - 1 = {x^2} - x - 1\] and after a simplification, I get this: \[{x^2} - 2x = 0\]

OpenStudy (anonymous):

That can be factored to x(x-2) = 0 Therefore x=0 and x=2 ?

OpenStudy (michele_laino):

correct!

OpenStudy (anonymous):

Thank you very much!

OpenStudy (michele_laino):

:)